OpenStudy (anonymous):

how do you graph this equation: y + 3 = -3(x + 2)

OpenStudy (mathteacher1729):

The SUPER IMPORTANT POINT to remember when graphing a line: >>>>> You need only TWO POINTS. >>>>> So let's take y + 3 = -3(x + 2) and set y = -3. That means 0 = -3(x+2) and that means x must equal -2. POINT ONE: (-2, -3) Now the slope of this line is -3/1 (go down three, go over one) So if we start from (-2,-3) and go down three, we land at (-2, -6) and if we go over one from that point we land at (-1, -6). POINT TWO: (-1,-6). Now plot those two points, and draw a line between them. :) Lemmie know if you have any questions about the process.

OpenStudy (anonymous):

i can't' go down -6 because the graph given only lets me go down to -3

OpenStudy (mathteacher1729):

Ah, ok. So here's what you do. "go down three and go over (right) one" also means "go UP three and go over (LEFT) one". Try that. :)

OpenStudy (anonymous):

ok i got it now, can you help me with this one? a deivery service charges a base price for an overnight delivery of a package plus an extra charge for each pound the package weighs. a customer is billed @22.85 for shipping a 3-pound package an $40 for shipping a 10-pound package, write an equation that gives the total cost of shipping a package as a function of the weight of the package